# Kitashirakawa Oiwakecho, Sakyo-ku, Kyoto 606-8502, Japan

### v1

@@ A.3 Proofs of (in)equivalence
  • Maximin $\Rightarrow$ Extremal: From Proposition A.2, it is sufficient to show that all maximin surfaces $\gamma_{m}(A;D)$ are extremal along $\Sigma_{m}(A;D)$. Consider a surface $\gamma$, $D$-homologous to $A$, and infinitesimally close to it. Definition A.1 implies that $\partial A=\partial\gamma$, and $\partial A^{c}=\partial A$ where $A^{c}$ is the complement of $A$ on $\Sigma|_{\mathcal{S}}$. Hence, $A^{c}\cup\gamma$ is a closed surface infinitesimally close to $\Sigma|_{\mathcal{S}}$, and it lies in on some Cauchy slice $\Sigma$. By Lemma A.4,
@@
  • Extremal $\Rightarrow$ Maximin is implied by Proposition A.2.
  • Extremal $\Leftrightarrow$ C-extremal: We proceed by contradiction and assume there exists a surface $\gamma_{\rm min}(A;D)$ $D$-homologous to $A$ with a positive measure subsystem on $\partial D$ while being the smallest area surface in $D$. By Lemma A.1, this surface is a minimin. Any infinitesimal deformation of $\gamma_{\rm min}(A;D)$ along the null direction directed towards the interior of $D$ must decrease the area by Definition 4.1. This is in contradiction with the assumption that $\gamma_{\rm min}(A;D)$ is a minimin.

### v3

@@ B.3 Proofs of (in)equivalence
  • Maximin $\Rightarrow$ Extremal: From Proposition B.2, it is sufficient to show that all maximin surfaces $\gamma_{m}(A;D)$ are extremal along $\Sigma_{m}(A;D)$. Consider a surface $\gamma$, $D$-homologous to $A$, and infinitesimally close to it. Definition B.1 implies that $\partial A=\partial\gamma$, and $\partial A^{c}=\partial A$ where $A^{c}$ is the complement of $A$ on $\Sigma|_{\mathcal{S}}$. Hence, $A^{c}\cup\gamma$ is a closed surface infinitesimally close to $\Sigma|_{\mathcal{S}}$, and it lies in on some Cauchy slice $\Sigma$. By Lemma B.4,
@@
  • Extremal $\Rightarrow$ Maximin is implied by Proposition B.2 and the above proof that all maximin surfaces are extremal along $\Sigma_{m}$.
  • Extremal $\Leftrightarrow$ C-extremal: We proceed by contradiction and assume there exists a surface $\gamma_{\rm min}(A;D)$ $D$-homologous to $A$ with a positive measure subsystem on $\partial D$ while being the smallest area surface in $D$. By Lemma B.1, this surface is a minimin. Any infinitesimal deformation of $\gamma_{\rm min}(A;D)$ along the null direction directed towards the interior of $D$ must decrease the area by Definition 4.2. This is in contradiction with the assumption that $\gamma_{\rm min}(A;D)$ is a minimin.